# Tuning LLMs with Contrastive Alignment Instructions for Machine Translation in Unseen, Low-resource Languages

### v1

@@ Abstract
Zhuoyuan Mao††thanks: Currently at Language Media Processing Lab., Kyoto University. Work done during Apple internship. and Yen Yu Apple kevinmzy@gmail.com, yen_yu@apple.com
@@ 1 Introduction
This study focused on machine translation (MT) to highlight the cross-lingual LLM adaptation challenge. The challenge lies in enabling translation for low-resource languages that often lack robust cross-lingual signals. We first explored the efficacy of fine-tuning LLMs with MT instructions (MTInstruct) in unseen, low-resource languages. MTInstruct is a method previously shown to bolster the translation proficiency of LLMs for supported languages Li et al. (2023). Subsequently, given that cross-lingual alignments are suboptimal in LLMs as a result of data scarcity of low-resource languages, we proposed contrastive alignment instructions (AlignInstruct) to explicitly provide cross-lingual supervision during MT fine-tuning. AlignInstruct is a cross-lingual discriminator formulated using statistical word alignments. Our approach was inspired by prior studies Lambert et al. (2012); Ren et al. (2019); Lin et al. (2020); Mao et al. (2022), which indicated the utility of word alignments in enhancing MT. In addition to AlignInstruct, we discussed two word-level cross-lingual instruction alternatives cast as generative tasks, for comparison with AlignInstruct.
Our experiments fine-tuned the BLOOMZ models Muennighoff et al. (2023) of varying sizes (1b1, 3b, and 7b1) for 24 unseen, low-resource languages, and evaluated translation on OPUS-100 Zhang et al. (2020) and Flores-200 Costa-jussà et al. (2022). We first showed that MTInstruct effectively induced the translation capabilities of LLMs for these languages. Building on the MTInstruct baseline, the multi-task learning combining AlignInstruct and MTInstruct resulted in stronger translation performance without the need for additional training corpora. The performance improved with larger BLOOMZ models, as illustrated in Fig. 1, indicating that AlignInstruct is particularly beneficial for larger LLMs during MT fine-tuning. When compared with the generative variants of AlignInstruct, our results indicated that discriminator-style instructions better complemented MTInstruct. Furthermore, merging AlignInstruct with its generative counterparts did not further improve translation quality, underscoring the efficacy and sufficiency of AlignInstruct in leveraging word alignments for MT.
In zero-shot translation evaluations on the OPUS benchmark, AlignInstruct exhibited improvements over the MTInstruct baseline in 30 zero-shot directions not involving English, when exclusively fine-tuned with three unseen languages (German, Dutch, and Russian). However, when the fine-tuning data incorporated supported languages (Arabic, French, and Chinese), the benefits of AlignInstruct were only evident in zero-shot translations where the target language was a supported language. In addition, to interpret the inherent modifications within the BLOOMZ models after applying MTInstruct or AlignInstruct, we conducted a visualization of the layer-wise cross-lingual alignment capabilities of the model representations.
@@ 2 Methodology
This section presents MTInstruct as the baseline, and AlignInstruct. The MTInstruct baseline involved fine-tuning LLMs using MT instructions. AlignInstruct dealt with the lack of cross-lingual signals stemming from the limited parallel training data in low-resource languages. The expectation was enhanced cross-lingual supervision cast as a discriminative task without extra training corpora. Following this, we introduced two generative variants of AlignInstruct for comparison.111We also discussed monolingual instructions for MT fine-tuning in App. D.2.
@@ 2.3.2 ReviseInstruct
ReviseInstruct was inspired by Ren et al. (2019) and Liu et al. (2020) for the notion of generating parallel words or phrases, thereby encouraging a model to encode cross-lingual alignments. A ReviseInstruct instruction contained a partially corrupted translation target, as well as a directive to identify and revise these erroneous tokens. Tokens are intentionally corrupted at the granularity of individual words, aligning with the word-level granularity in AlignInstruct and HintInstruct. ReviseInstruct follows the instruction pattern:333We illustrated examples of HintInstruct and ReviseInstruct in App. D.1 for reference.
@@ 3.1 Backbone Models and Unseen Languages
BLOOMZ+24 Tuning BLOOMZ-7b1, BLOOMZ-3b, and BLOOMZ-1b1444https://huggingface.co/bigscience/bloomz for unseen, low-resource languages. These experiments aimed to: (1) assess the effectiveness of AlignInstruct in multilingual, low-resource scenarios; (2) offer comparison across various model sizes. We used the OPUS-100 Zhang et al. (2020)555https://opus.nlpl.eu/opus-100.php datasets as training data. OPUS-100 is an English-centric parallel corpora, with around M parallel sentences in total for 24 selected languages, averaging k sentence pairs for each language and English. Refer to App. A for training data statistics. We used OPUS-100 and Flores-200 Costa-jussà et al. (2022)666https://github.com/facebookresearch/flores/blob/main/flores200/README.md for evaluating translation between English and unseen languages (48 directions in total) on in-domain and out-of-domain test sets, respectively. The identical prompt as introduced in Sec. 2.1 was employed for inference. Inferences using alternative MT prompts are discussed in App.E.
@@ 3.2 Training Details and Curricula
Mixed Fine-tuning Chu et al. (2017) arranged the two aforementioned curricula to start with MT+Align, followed by MTInstruct, denoted as MT+AlignMT.
@@ 4.1 BLOOMZ+24 Results
Tab. 1 shows the scores for the unmodified BLOOMZ models, as well as BLOOMZ+24 under MTInstruct, AlignInstruct, and the three distinct curricula. Non-trivial improvements in all metrics were evident for BLOOMZ+24 under MTInstruct. This suggests that MTInstruct can induce translation capabilities in unseen languages. Applying AlignInstruct and MTInstruct via the curricula further showed better scores than the baselines, suggesting the role of AlignInstruct as complementing MTInstruct. AlignMT was an exception, performing similarly to MTInstruct. This may indicate AlignInstruct’s complementarity depends on its cadence relative to MTInstruct in a curriculum.
Superior OPUS and Flores scores under the xxen direction were evident, compared to the reverse direction, enxx. This suggests that our treatments induced understanding capabilities more than generative ones. This may be attributed to the fact that BLOOMZ had significant exposure to English, and that we used English-centric corpora. Finally, we noted the inferior performance of Flores than OPUS. This speaks to the challenge of instilling translation abilities in unseen languages when dealing with the out-of-domain MT task. Our future work will focus on enhancing the domain generalization capabilities of LLM fine-tuning in MT tasks.
@@ 4.2 Assessing AlignInstruct Variants
From the results reported in Tab. 3, we observed the objectives with AlignInstruct consistently outperformed those with HintInstruct or ReviseInstruct across metrics and model sizes. Namely, easy, discriminative instructions, rather than hard, generative ones, may be preferred for experiments under similar data constraints. The low-resource constraint likely made MTInstruct more sensitive to the difficulty of its accompanying tasks.
@@ 4.3 BLOOMZ+3 Zero-shot Evaluation
Tab. 5 reports the results of the two settings, de-nl-ru and ar-de-fr-nl-ru-zh. Results of MT+Align+Hint+Revise and pivot-based translation are reported in App. C and F. In the de-nl-ru setting, where BLOOMZ was fine-tuned with the three unseen languages, we noticed MT+Align consistently outperformed the MTInstruct baseline across all evaluated zero-shot directions. Notably, MT+Align enhanced the translation quality for unseenseen and seenunseen directions compared to w/o fine-tuning and MTInstruct, given that the model was solely fine-tuned on de, nl, and ru data. This suggested AlignInstruct not only benefits the languages supplied in the data but also has a positive impact on other languages through cross-lingual alignment supervision. In terms of supervised directions involving English, we noticed performance improvements associated with unseen languages, and regression in seen ones. The regression may be attributed to forgetting for the absence of seen languages in fine-tuning data. Indeed, continuous exposure to English maintained the translation quality for seenen. As LoRA is modular, the regression can be mitigated by detaching the LoRA parameters for seen languages.
@@ 5 Related Work
Prompting LLMs for MT LLMs have shown good performance for multilingual MT through few-shot in-context learning (ICL) Jiao et al. (2023). Vilar et al. (2023) showed that high-quality examples can improve MT based on PaLM Chowdhery et al. (2022). Agrawal et al. (2023) and Zhang et al. (2023a) explored strategies to compose better examples for few-shot prompting for XGLM-7.5B Lin et al. (2022) and GLM-130B Zeng et al. (2023). Ghazvininejad et al. (2023), Peng et al. (2023), and Moslem et al. (2023) claimed that dictionary-based hints and domain-specific style information can improve prompting OPT Zhang et al. (2022), GPT-3.5 Brown et al. (2020), and BLOOM Scao et al. (2022) for MT. He et al. (2023) used LLMs to mine useful knowledge for prompting GPT-3.5 for MT.
@@ 6 Conclusion
PEFT Methods and Adapters As discussed in the BLOOM+1 paper Yong et al. (2023), alternative PEFT techniques, such as (IA) Liu et al. (2022), have the potential to enhance the adaptation performance of LLM pre-training for previously unseen languages. These approaches are worth exploring for MT fine-tuning in such languages, in addition to the LoRA methods employed in this study. Furthermore, our exploration was limited to fine-tuning multiple languages using shared additional parameters. Investigating efficient adaptation through the use of the mixture of experts (MoE) approach for MT tasks Fan et al. (2021); Costa-jussà et al. (2022); Mohammadshahi et al. (2022); Koishekenov et al. (2023); Xu et al. (2023b) presents another intriguing avenue for LLM fine-tuning.
@@ Appendix D Representation Change of BLOOMZ+3
D.1 Examples of HintInstruct and ReviseInstruct
@@
D.2 Assessing Monolingual Instructions
@@
Appendix E Inference using Different MT Prompts
@@
Appendix F Zero-shot Translation using English as Pivot
@@
Appendix G Per Language Result Details of BLOOMZ+24 and BLOOMZ+3
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET |
| ar-de | 1.4 | 14.8 | 56.19 | en-ar | 11.1 | 32.4 | 75.66 |
| ar-fr | 21.9 | 46.1 | 74.19 | en-de | 12.2 | 29.2 | 59.16 |
| ar-nl | 0.6 | 11.2 | 56.59 | en-fr | 26.8 | 49.2 | 77.42 |
| ar-ru | 3.1 | 6.2 | 48.41 | en-nl | 2.0 | 16.0 | 46.52 |
| ar-zh | 18.4 | 14.4 | 73.65 | en-ru | 5.7 | 16.1 | 49.00 |
| de-ar | 2.0 | 17.8 | 64.91 | en-zh | 22.5 | 17.0 | 77.90 |
| de-fr | 12.0 | 33.4 | 63.45 | avg. | 13.38 | 26.65 | 64.28 |
| de-nl | 3.7 | 17.9 | 47.30 | ||||
| de-ru | 1.3 | 11.8 | 45.53 | ||||
| de-zh | 8.9 | 7.6 | 61.52 | ||||
| fr-ar | 11.2 | 33.4 | 74.20 | BLEU | chrF++ | COMET | |
| fr-de | 4.6 | 23.4 | 48.83 | ar-en | 26.7 | 48.4 | 78.12 |
| fr-nl | 2.8 | 17.2 | 52.14 | de-en | 21.1 | 38.5 | 71.99 |
| fr-ru | 3.1 | 10.4 | 45.12 | fr-en | 27.7 | 49.8 | 79.46 |
| fr-zh | 20.9 | 17.0 | 76.20 | nl-en | 12.3 | 31.1 | 61.29 |
| nl-ar | 1.3 | 13.2 | 59.46 | ru-en | 17.9 | 36.6 | 68.40 |
| nl-de | 5.9 | 22.8 | 46.49 | zh-en | 24.5 | 47.9 | 77.08 |
| nl-fr | 9.6 | 29.6 | 58.30 | avg. | 21.70 | 42.05 | 72.72 |
| nl-ru | 0.8 | 9.0 | 42.83 | ||||
| nl-zh | 3.3 | 3.7 | 53.96 | ||||
| ru-ar | 6.5 | 25.3 | 68.38 | ||||
| ru-de | 2.0 | 17.0 | 48.06 | ||||
| ru-fr | 15.7 | 38.7 | 67.54 | ||||
| ru-nl | 0.5 | 10.5 | 46.14 | ||||
| ru-zh | 10.7 | 11.3 | 67.18 | ||||
| zh-ar | 8.6 | 29.7 | 73.47 | ||||
| zh-de | 1.6 | 17.6 | 49.90 | ||||
| zh-fr | 20.7 | 44.1 | 75.79 | ||||
| zh-nl | 0.6 | 10.4 | 48.53 | ||||
| zh-ru | 2.9 | 8.6 | 44.13 | ||||
| avg. | 6.89 | 19.14 | 57.95 | ||||
| seenseen | 16.95 | 30.78 | 74.58 | enseen | 20.13 | 32.87 | 76.99 |
| seenunseen | 2.30 | 13.31 | 49.98 | enunseen | 6.63 | 20.43 | 51.56 |
| unseenseen | 7.78 | 20.07 | 62.74 | seenen | 26.30 | 48.70 | 78.22 |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET |
| ar-de | 4.7 | 20.9 | 56.43 | en-ar | 9.1 | 27.2 | 71.47 |
| ar-fr | 20.8 | 42.5 | 71.47 | en-de | 19.8 | 36.1 | 66.53 |
| ar-nl | 7.2 | 22.9 | 58.29 | en-fr | 23.0 | 44.5 | 74.98 |
| ar-ru | 5.0 | 21.0 | 54.73 | en-nl | 15.5 | 36.1 | 64.76 |
| ar-zh | 14.0 | 12.4 | 67.94 | en-ru | 14.2 | 30.3 | 62.82 |
| de-ar | 2.4 | 16.2 | 64.53 | en-zh | 20.7 | 17.9 | 74.97 |
| de-fr | 11.9 | 31.2 | 64.44 | avg. | 17.05 | 32.02 | 69.26 |
| de-nl | 9.4 | 28.1 | 54.22 | ||||
| de-ru | 5.1 | 19.6 | 55.41 | ||||
| de-zh | 4.2 | 5.8 | 55.26 | ||||
| fr-ar | 10.1 | 29.1 | 70.72 | BLEU | chrF++ | COMET | |
| fr-de | 8.6 | 27.7 | 53.77 | ar-en | 26.5 | 46.9 | 76.92 |
| fr-nl | 10.3 | 30.1 | 57.55 | de-en | 27.0 | 44.0 | 76.97 |
| fr-ru | 7.9 | 26.0 | 56.82 | fr-en | 27.5 | 49.0 | 78.80 |
| fr-zh | 18.1 | 18.5 | 72.24 | nl-en | 21.8 | 41.3 | 73.99 |
| nl-ar | 2.0 | 15.1 | 63.73 | ru-en | 24.8 | 43.6 | 74.23 |
| nl-de | 9.7 | 28.1 | 52.58 | zh-en | 23.2 | 45.3 | 76.83 |
| nl-fr | 13.2 | 32.3 | 65.17 | avg. | 25.13 | 45.02 | 76.29 |
| nl-ru | 5.1 | 18.6 | 55.13 | ||||
| nl-zh | 3.0 | 5.4 | 54.34 | ||||
| ru-ar | 5.9 | 15.0 | 60.36 | ||||
| ru-de | 5.6 | 23.8 | 52.66 | ||||
| ru-fr | 17.9 | 38.4 | 68.66 | ||||
| ru-nl | 6.2 | 22.5 | 54.41 | ||||
| ru-zh | 7.5 | 13.6 | 61.40 | ||||
| zh-ar | 6.7 | 22.1 | 67.48 | ||||
| zh-de | 3.3 | 19.6 | 51.75 | ||||
| zh-fr | 17.4 | 38.9 | 73.00 | ||||
| zh-nl | 4.8 | 19.3 | 54.41 | ||||
| zh-ru | 3.5 | 17.9 | 49.02 | ||||
| avg. | 8.38 | 22.75 | 59.93 | ||||
| seenseen | 14.52 | 27.25 | 70.48 | enseen | 17.60 | 29.87 | 73.81 |
| seenunseen | 6.14 | 22.82 | 54.75 | enunseen | 16.50 | 34.17 | 64.70 |
| unseenseen | 7.56 | 19.22 | 61.99 | seenen | 25.73 | 47.07 | 77.52 |

### v2

@@ Abstract
Zhuoyuan Mao††thanks: Currently at Sony Group Corporation. Work done during Apple internship. and Yen Yu Apple kevinmzy@gmail.com, yen_yu@apple.com
@@ 1 Introduction
This study focused on machine translation (MT) to highlight the cross-lingual LLM adaptation challenge. The challenge lies in enabling translation for low-resource languages that often lack robust cross-lingual signals. We first explored the efficacy of fine-tuning LLMs with MT instructions (MTInstruct) in unseen, low-resource languages. MTInstruct is a method previously shown to bolster the translation proficiency of LLMs for supported languages Li et al. (2023). Subsequently, given that cross-lingual alignments are suboptimal in LLMs as a result of data scarcity of low-resource languages, we proposed contrastive alignment instructions (AlignInstruct) to explicitly provide cross-lingual supervision during MT fine-tuning. AlignInstruct is a cross-lingual discriminator formulated using statistical word alignments. Our approach was inspired by prior studies Lambert et al. (2012); Ren et al. (2019); Lin et al. (2020); Mao et al. (2022), which indicated the utility of word alignments in enhancing MT. In addition to AlignInstruct, we discussed two word-level cross-lingual instruction alternatives cast as generative tasks for comparison with AlignInstruct.
Our experiments fine-tuned the BLOOMZ models Muennighoff et al. (2023) of varying sizes (1b1, 3b, and 7b1) for 24 unseen, low-resource languages, and evaluated translation on OPUS-100 Zhang et al. (2020) and Flores-200 Costa-jussà et al. (2022). We first showed that MTInstruct effectively induced the translation capabilities of LLMs for these languages. Building on the MTInstruct baseline, the multi-task learning combining AlignInstruct and MTInstruct resulted in stronger translation performance without the need for additional training corpora. The performance improved with larger BLOOMZ models, as illustrated in Fig. 1, indicating that AlignInstruct is particularly beneficial for larger LLMs during MT fine-tuning. When compared with the generative variants of AlignInstruct, our results indicated that discriminative instructions better complemented MTInstruct. Furthermore, merging AlignInstruct with its generative counterparts did not further improve translation quality, underscoring the efficacy and sufficiency of AlignInstruct in leveraging word alignments for MT.
In zero-shot translation evaluation on the OPUS benchmark, AlignInstruct exhibited improvements over the MTInstruct baseline in 30 zero-shot directions between non-English languages, when exclusively fine-tuned with three unseen languages (German, Dutch, and Russian). However, when incorporating supported languages (Arabic, French, and Chinese) the benefits of AlignInstruct were only evident in zero-shot translations where the target language was a supported language. In addition, to interpret the inherent modifications within the BLOOMZ models after applying MTInstruct or AlignInstruct, we conducted a visualization of the layer-wise cross-lingual alignment capabilities of the model representations.
@@ 2 Methodology
This section presents MTInstruct as the baseline, and AlignInstruct. The MTInstruct baseline involved fine-tuning LLMs using MT instructions. AlignInstruct dealt with the lack of cross-lingual signals stemming from the limited parallel training data in low-resource languages. The expectation was enhanced cross-lingual supervision cast as a discriminative task without extra training corpora. Following this, we introduced two generative variants of AlignInstruct for comparison.111We also discussed monolingual instructions for MT fine-tuning in App. F.
@@ 2.3.2 ReviseInstruct
ReviseInstruct was inspired by Ren et al. (2019) and Liu et al. (2020) for the notion of generating parallel words or phrases, thereby encouraging a model to encode cross-lingual alignments. A ReviseInstruct instruction contained a partially corrupted translation target, as well as a directive to identify and revise these erroneous tokens. Tokens are intentionally corrupted at the granularity of individual words, aligning with the word-level granularity in AlignInstruct and HintInstruct. ReviseInstruct follows the instruction pattern:333We illustrated examples of HintInstruct and ReviseInstruct in App. E for reference.
@@ 3.1 Backbone Models and Unseen Languages
BLOOMZ+24 Tuning BLOOMZ-7b1, BLOOMZ-3b, and BLOOMZ-1b1444https://huggingface.co/bigscience/bloomz for unseen, low-resource languages. These experiments aimed to: (1) assess the effectiveness of AlignInstruct in multilingual, low-resource scenarios; (2) offer comparison across various model sizes. We used the OPUS-100 Zhang et al. (2020)555https://opus.nlpl.eu/opus-100.php datasets as training data. OPUS-100 is an English-centric parallel corpora, with around M parallel sentences in total for 24 selected languages, averaging k sentence pairs for each language and English. Refer to App. A for training data statistics. We used OPUS-100 and Flores-200 Costa-jussà et al. (2022)666https://github.com/facebookresearch/flores/blob/main/flores200/README.md for evaluating translation between English and unseen languages (48 directions in total) on in-domain and out-of-domain test sets, respectively. The identical prompt as introduced in Sec. 2.1 was employed for inference. Inferences using alternative MT prompts are discussed in App. G.
@@ 3.2 Training Details and Curricula
Mixed Fine-tuning Chu et al. (2017) arranged the above curricula to start with MT+Align, followed by MTInstruct, denoted as MT+AlignMT.
@@ 4.1 BLOOMZ+24 Results
Tab. 1 shows the scores for the unmodified BLOOMZ models, as well as BLOOMZ+24 under MTInstruct, AlignInstruct, and the three distinct curricula. Non-trivial improvements in all metrics were evident for BLOOMZ+24 under MTInstruct. This suggests that MTInstruct can induce translation capabilities in unseen languages. Applying AlignInstruct and MTInstruct via the curricula further showed better scores than the baselines, suggesting the role of AlignInstruct as complementing MTInstruct. AlignMT was an exception, performing similarly to MTInstruct. This may indicate the effect of AlignInstruct depends on its cadence relative to MTInstruct in a curriculum.
Superior OPUS and Flores scores under the xxen direction were evident, compared to the reverse direction, enxx. This suggests that our treatments induced understanding capabilities more than generative ones. This may be attributed to the fact that BLOOMZ had significant exposure to English, and that we used English-centric corpora. Finally, we noted the inferior performance of Flores than OPUS. This speaks to the challenge of instilling out-of-domain translation abilities in unseen languages. Our future work will focus on enhancing the domain generalization capabilities of LLM fine-tuning in MT tasks.
@@ 4.2 Assessing AlignInstruct Variants
From Tab. 3, we observed the objectives with AlignInstruct consistently outperformed those with HintInstruct or ReviseInstruct across metrics and model sizes. Namely, easy, discriminative instructions, rather than hard, generative ones, may be preferred for experiments under similar data constraints. The low-resource constraint likely made MTInstruct more sensitive to the difficulty of its accompanying tasks.
@@ 4.3 BLOOMZ+3 Zero-shot Evaluation
Tab. 5 reports the results of the two settings, de-nl-ru and ar-de-fr-nl-ru-zh. Results of MT+Align+Hint+Revise and pivot-based translation are reported in App. C and H. In the de-nl-ru setting, where BLOOMZ was fine-tuned with the three unseen languages, we noticed MT+Align consistently outperformed the MTInstruct baseline across all evaluated zero-shot directions. Notably, MT+Align enhanced the translation quality for unseenseen and seenunseen directions compared to w/o fine-tuning and MTInstruct, given that the model was solely fine-tuned on de, nl, and ru data. This suggested AlignInstruct not only benefits the languages supplied in the data but also has a positive impact on other languages through cross-lingual alignment supervision. In terms of supervised directions involving English, we noticed performance improvements associated with unseen languages, and regression in seen ones. The regression may be attributed to forgetting for the absence of seen languages in fine-tuning data. Indeed, continuous exposure to English maintained the translation quality for seenen. As LoRA is modular, the regression can be mitigated by detaching the LoRA parameters for seen languages.
@@ 5 Related Work
Prompting LLMs for MT LLMs have shown good performance for multilingual MT through few-shot in-context learning (ICL) Jiao et al. (2023). Agrawal et al. (2023) and Zhang et al. (2023a) explored strategies to compose better examples for ICL for XGLM-7.5B Lin et al. (2022) and GLM-130B Zeng et al. (2023). Ghazvininejad et al. (2023), Peng et al. (2023), and Moslem et al. (2023) claimed that dictionary-based hints and domain-specific style information can improve prompting OPT Zhang et al. (2022), GPT-3.5 Brown et al. (2020), and BLOOM Scao et al. (2022) for MT. He et al. (2023) used LLMs to mine useful knowledge for prompting GPT-3.5 for MT.
@@ 6 Conclusion
PEFT Methods and Adapters As discussed in the BLOOM+1 paper Yong et al. (2023), alternative PEFT techniques, such as (IA)3 Liu et al. (2022), have the potential to enhance the adaptation performance of LLM pre-training for previously unseen languages. These approaches are worth exploring for MT fine-tuning in such languages, in addition to the LoRA methods employed in this study. Furthermore, our exploration was limited to fine-tuning multiple languages using shared additional parameters. Investigating efficient adaptation through the use of the mixture of experts (MoE) approach for MT tasks Fan et al. (2021); Costa-jussà et al. (2022); Mohammadshahi et al. (2022); Koishekenov et al. (2023); Xu et al. (2023b) presents another intriguing avenue for LLM fine-tuning.
@@ Appendix D Representation Change of BLOOMZ+3
Appendix E Examples of HintInstruct and ReviseInstruct
@@
Appendix F Assessing Monolingual Instructions
@@
Appendix G Inference using Different MT Prompts
@@
Appendix H Zero-shot Translation using English as Pivot
@@
Appendix I Per Language Result Details of BLOOMZ+24 and BLOOMZ+3
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 1.4 | 14.8 | 56.19 | en-ar | 11.1 | 32.4 | 75.66 | |
| ar-fr | 21.9 | 46.1 | 74.19 | en-de | 12.2 | 29.2 | 59.16 | |
| ar-nl | 0.6 | 11.2 | 56.59 | en-fr | 26.8 | 49.2 | 77.42 | |
| ar-ru | 3.1 | 6.2 | 48.41 | en-nl | 2.0 | 16.0 | 46.52 | |
| ar-zh | 18.4 | 14.4 | 73.65 | en-ru | 5.7 | 16.1 | 49.00 | |
| de-ar | 2.0 | 17.8 | 64.91 | en-zh | 22.5 | 17.0 | 77.90 | |
| de-fr | 12.0 | 33.4 | 63.45 | avg. | 13.38 | 26.65 | 64.28 | |
| de-nl | 3.7 | 17.9 | 47.30 | |||||
| de-ru | 1.3 | 11.8 | 45.53 | |||||
| de-zh | 8.9 | 7.6 | 61.52 | |||||
| fr-ar | 11.2 | 33.4 | 74.20 | BLEU | chrF++ | COMET | ||
| fr-de | 4.6 | 23.4 | 48.83 | ar-en | 26.7 | 48.4 | 78.12 | |
| fr-nl | 2.8 | 17.2 | 52.14 | de-en | 21.1 | 38.5 | 71.99 | |
| fr-ru | 3.1 | 10.4 | 45.12 | fr-en | 27.7 | 49.8 | 79.46 | |
| fr-zh | 20.9 | 17.0 | 76.20 | nl-en | 12.3 | 31.1 | 61.29 | |
| nl-ar | 1.3 | 13.2 | 59.46 | ru-en | 17.9 | 36.6 | 68.40 | |
| nl-de | 5.9 | 22.8 | 46.49 | zh-en | 24.5 | 47.9 | 77.08 | |
| nl-fr | 9.6 | 29.6 | 58.30 | avg. | 21.70 | 42.05 | 72.72 | |
| nl-ru | 0.8 | 9.0 | 42.83 | |||||
| nl-zh | 3.3 | 3.7 | 53.96 | |||||
| ru-ar | 6.5 | 25.3 | 68.38 | |||||
| ru-de | 2.0 | 17.0 | 48.06 | |||||
| ru-fr | 15.7 | 38.7 | 67.54 | |||||
| ru-nl | 0.5 | 10.5 | 46.14 | |||||
| ru-zh | 10.7 | 11.3 | 67.18 | |||||
| zh-ar | 8.6 | 29.7 | 73.47 | |||||
| zh-de | 1.6 | 17.6 | 49.90 | |||||
| zh-fr | 20.7 | 44.1 | 75.79 | |||||
| zh-nl | 0.6 | 10.4 | 48.53 | |||||
| zh-ru | 2.9 | 8.6 | 44.13 | |||||
| avg. | 6.89 | 19.14 | 57.95 | |||||
| seenseen | 16.95 | 30.78 | 74.58 | enseen | 20.13 | 32.87 | 76.99 | |
| seenunseen | 2.30 | 13.31 | 49.98 | enunseen | 6.63 | 20.43 | 51.56 | |
| unseenseen | 7.78 | 20.07 | 62.74 | seenen | 26.30 | 48.70 | 78.22 | |
@@
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 4.7 | 20.9 | 56.43 | en-ar | 9.1 | 27.2 | 71.47 | |
| ar-fr | 20.8 | 42.5 | 71.47 | en-de | 19.8 | 36.1 | 66.53 | |
| ar-nl | 7.2 | 22.9 | 58.29 | en-fr | 23.0 | 44.5 | 74.98 | |
| ar-ru | 5.0 | 21.0 | 54.73 | en-nl | 15.5 | 36.1 | 64.76 | |
| ar-zh | 14.0 | 12.4 | 67.94 | en-ru | 14.2 | 30.3 | 62.82 | |
| de-ar | 2.4 | 16.2 | 64.53 | en-zh | 20.7 | 17.9 | 74.97 | |
| de-fr | 11.9 | 31.2 | 64.44 | avg. | 17.05 | 32.02 | 69.26 | |
| de-nl | 9.4 | 28.1 | 54.22 | |||||
| de-ru | 5.1 | 19.6 | 55.41 | |||||
| de-zh | 4.2 | 5.8 | 55.26 | |||||
| fr-ar | 10.1 | 29.1 | 70.72 | BLEU | chrF++ | COMET | ||
| fr-de | 8.6 | 27.7 | 53.77 | ar-en | 26.5 | 46.9 | 76.92 | |
| fr-nl | 10.3 | 30.1 | 57.55 | de-en | 27.0 | 44.0 | 76.97 | |
| fr-ru | 7.9 | 26.0 | 56.82 | fr-en | 27.5 | 49.0 | 78.80 | |
| fr-zh | 18.1 | 18.5 | 72.24 | nl-en | 21.8 | 41.3 | 73.99 | |
| nl-ar | 2.0 | 15.1 | 63.73 | ru-en | 24.8 | 43.6 | 74.23 | |
| nl-de | 9.7 | 28.1 | 52.58 | zh-en | 23.2 | 45.3 | 76.83 | |
| nl-fr | 13.2 | 32.3 | 65.17 | avg. | 25.13 | 45.02 | 76.29 | |
| nl-ru | 5.1 | 18.6 | 55.13 | |||||
| nl-zh | 3.0 | 5.4 | 54.34 | |||||
| ru-ar | 5.9 | 15.0 | 60.36 | |||||
| ru-de | 5.6 | 23.8 | 52.66 | |||||
| ru-fr | 17.9 | 38.4 | 68.66 | |||||
| ru-nl | 6.2 | 22.5 | 54.41 | |||||
| ru-zh | 7.5 | 13.6 | 61.40 | |||||
| zh-ar | 6.7 | 22.1 | 67.48 | |||||
| zh-de | 3.3 | 19.6 | 51.75 | |||||
| zh-fr | 17.4 | 38.9 | 73.00 | |||||
| zh-nl | 4.8 | 19.3 | 54.41 | |||||
| zh-ru | 3.5 | 17.9 | 49.02 | |||||
| avg. | 8.38 | 22.75 | 59.93 | |||||
| seenseen | 14.52 | 27.25 | 70.48 | enseen | 17.60 | 29.87 | 73.81 | |
| seenunseen | 6.14 | 22.82 | 54.75 | enunseen | 16.50 | 34.17 | 64.70 | |
| unseenseen | 7.56 | 19.22 | 61.99 | seenen | 25.73 | 47.07 | 77.52 | |